# FuncGenFoil: Airfoil Generation and Editing Model in Function Space

### v1

@@ 3.2 Airfoil Editing
The airfoil editing task enables the user to modify parts of the geometry of a given airfoil, effectively generating a new airfoil geometry while preserving the user-edited sections, denoted as $u_{1}^{\delta}$. We achieve this by an optimization method that maximum a posteriori (MAP) estimation, $\max p(u_{1}^{{}^{\prime}}\mid u_{1}^{\delta})$ where $p(u_{1}^{{}^{\prime}}\mid u_{1}^{\delta})$ is a probabilistic model that constrains the optimized airfoil $u_{1}^{{}^{\prime}}$ fulfilling the editing requirements and following the generation prior.
@@
$\max_{u_{1}^{{}^{\prime}}}p(u_{1}^{{}^{\prime}}\mid u_{1}^{\delta})\;=\;\frac{p(u_{1}^{\delta}\mid u_{1}^{{}^{\prime}})\cdot p(u_{1}^{{}^{\prime}})}{p(u_{1}^{\delta})}$ (6)
$\Rightarrow$ $\max_{u_{1}^{{}^{\prime}}}\log{p(u_{1}^{\delta}|u_{1}^{{}^{\prime}})}+\log{p(u_{1}^{{}^{\prime}})}-\log{p(u_{1}^{\delta})},$
where $p(u_{1}^{\delta}\mid u_{1}^{{}^{\prime}})$ is a Gaussian measure, so its log term becomes a Mean Square Error (MSE) between $u_{1}^{\delta}$ and $u_{1}^{{}^{\prime}}$, constraining the user edited parts in $u_{1}^{{}^{\prime}}$ keeping consistent with $u_{1}^{\delta}$. $p(u_{1}^{{}^{\prime}})$ is the prior supported by the trained generative model so its log value is derived by Hutchinson trace estimator Hutchinson (1990); Grathwohl et al. (2019) effectively. $p(u_{1}^{\delta})$ is the marginal likelihood, which does not depend on $u_{1}^{{}^{\prime}}$. The final optimization target could be written as:
@@
where $\sigma$ is noise, $\Delta$ denotes point indices in edited part $u_{1}^{\delta}$.
For more realistic generation results, the optimization does not directly adjust $u_{1}^{{}^{\prime}}$; instead, we optimize $u_{1}^{{}^{\prime}}$ indirectly by fine-tuning the entire generative model for a few iterations. The fine-tuning process is illustrated in Fig. 4, while its details are provided in Alg. 1 in the main text. Specifically, we first initialize $u^{\prime}_{1}$ as a resample data of $u_{1}$, by extracting its latent code, denoted as $\alpha_{\theta}$, via the inverse of our generative model and re-generating $u_{1}^{{}^{\prime}}$ from this code. Then we treat Equation 7 as the loss function to train $\theta$ of the velocity operator and $\alpha_{\theta}$ simultaneously. After the model training, the new $u_{1}^{{}^{\prime}}$ is the new edited generation results.
Algorithm 1 Airfoil Editing.
Output: new airfoil $\{y_{i}\}$ at resolution $d$.
2:  Set $\{a_{\theta,i}\}=\{u_{0,i}\}$.
3:  while finetuning… do
4:     Sample $\{u_{1,i}^{{}^{\prime}}\}$ through $v_{\theta}$ from $\{a_{\theta,i}\}$.
5:     Compute $\mathcal{L}_{\mathrm{MAP}}$.
6:     Compute gradient and update $\theta$.
7:  end while
8:  Sample $\{u_{1,i}^{{}^{\prime}}\}$ through $v_{\theta}$ from $\{a_{\theta,i}\}$.
9:  return $\{y_{i}\}=\{u_{1,i}^{{}^{\prime}}\}$
@@ 4.1 Experiment Settings
Tasks. We evaluate two tasks in the airfoil inverse design problem: conditional airfoil generation and freestyle airfoil editing. In the conditional generation task, the model is given a set of 11 geometric parameters that describe the airfoil geometry. Detailed parameter definitions are provided in the Supplementary Materials. The model must generate airfoils that satisfy these geometric constraints. In the freestyle editing task, the model takes an original airfoil and a target modification, such as adjusting the position of a specific point on the airfoil. The selected point can be any location on the airfoil surface. The model must generate an airfoil that reflects the specified modification.
Metrics. We adopt the metrics introduced in AFBench Liu et al. (2024) to evaluate the generated airfoils:
@@
Label Error measures the difference between the physical parameters of the generated or edited airfoil and the intended target parameters, calculated as $\sigma_{i}=\left|\hat{p}_{i}-p_{i}\right|$, where $\sigma_{i}$ is the label error for the $i$-th parameter, $\hat{p}_{i}$ is the $i$-th geometric parameter from the generated airfoil, and $p_{i}$ is the corresponding target parameter. Smaller values indicate better alignment with the target parameters.
@@
Diversity quantifies the variety of generated airfoils, calculated as $D=\frac{1}{n}\sum_{i=1}^{n}\log\det(L_{S_{i}})$, where $n$ is the number of samples, and $L_{S_{i}}$ is the similarity matrix of the $i$-th subset, computed based on Euclidean distances between airfoils in the subset. Higher values indicate greater diversity among generated airfoils.
Smoothness measures the geometric smoothness of the generated airfoils, calculated as:
$\small M=\sum_{i=1}^{N}\text{Distance}(P_{n},P_{n-1}P_{n+1}),$ (8)
@@
where $P_{n}$ is the $n$-th point, and $P_{n-1}P_{n+1}$ is the line segment between adjacent points. The function $\text{Distance}(P_{n},P_{n-1}P_{n+1})$ computes the perpendicular distance from $P_{n}$ to this line segment. Smaller values indicate better geometric quality.
@@
Datasets. To benchmark our method, we conduct experiments on three datasets: UIUC Selig (1996), Supercritical Airfoil (Super), and AF-200K. UIUC contains 1,600 designed airfoil geometries. Super focuses on supercritical airfoils and includes approximately 20,000 airfoil samples. AF-200K consists of 200,000 highly diversified airfoil samples.
@@
Baselines. We include the baseline models proposed in AFBench Liu et al. (2024), specifically the conditional VAE (CVAE), conditional GAN (CGAN), the modified VAE with PARSEC parameters and control keypoints (PK-VAE), along with PK-GAN, PKVAE-GAN, the U-Net-based PK-DIFF, and the transformer-based PK-DIT.
@@
Table 1: Quantitative comparison between FuncGenFoil and baseline methods across different datasets for conditional generation task. Label error, diversity, and smoothness of the generated airfoils are reported.
@@
$\sigma_{1}$ $\sigma_{2}$ $\sigma_{3}$ $\sigma_{4}$ $\sigma_{5}$ $\sigma_{6}$ $\sigma_{7}$ $\sigma_{8}$ $\sigma_{9}$ $\sigma_{10}$ $\sigma_{11}$ $\bar{\sigma}$
CVAE AF-200K 72.9 52.5 35.2 15900 99 95 29000 19.1 15.3 46 104 4131 -155.4 7.09
CGAN AF-200K 107 85.0 54.4 23200 143 137 59600 25.3 22.3 53 129 7596 -120.5 7.31
PK-VAE AF-200K 63.0 47.9 31.3 8620 66 64 17100 13.5 9.3 33 78 2375 -150.1 5.93
PK-GAN AF-200K 81.8 63.0 47.0 21030 120 117 32470 22.5 19.6 50 122 4923 -112.3 3.98
PKVAE-GAN AF-200K 56.8 31.7 31.0 5650 46 43 12000 9.1 5.1 28 63 1633 -129.6 2.89
PK-DIT AF-200K 11.2 32.3 15.4 1050 13 11.5 9790 0.5 0.5 23 24 997 -93.2 1.04
FuncGenFoil AF-200K 1.84 17.4 0.56 721 47.7 0.98 1676 0.45 0.65 160 174 255 -71.6 1.41
PK-DIT UIUC 63.8 51.4 33.6 11830 87 84.9 25700 16.9 11.9 36 98 3456 -141.7 6.03
FuncGenFoil UIUC 14.2 12.7 0.86 434.9 90.7 3.7 3559 0.55 0.69 129.9 135.2 398.4 -66.3 1.17
PK-DIT Super 52.0 35.0 24.0 3010 29 33.2 10500 8.3 2.6 27 33 1250 -123.4 1.97
FuncGenFoil Super 0.73 1.9 0.09 53.2 1.07 0.07 56.4 0.06 0.06 10.5 13.6 12.5 -108.8 0.98
Table 2: Quantitative evaluation of the airfoil editing task across different edit scales. The mean squared error (MSE) between the generated and target airfoils as well as the smoothness of the generated airfoils are reported.
@@
Table 3: Quantitative evaluation of our model (training resolution=257) across different sampling resolutions for conditional generation task.
[…]
$\sigma_{1}$ $\sigma_{2}$ $\sigma_{3}$ $\sigma_{4}$ $\sigma_{5}$ $\sigma_{6}$ $\sigma_{7}$ $\sigma_{8}$ $\sigma_{9}$ $\sigma_{10}$ $\sigma_{11}$ $\bar{\sigma}$
Super 257 0.73 1.9 0.09 53.2 1.07 0.07 56.4 0.06 0.06 10.5 13.6 12.5 -108.8 0.98
513 0.74 2.59 0.13 55.1 1.54 0.09 57.7 0.07 0.08 41.1 40.0 18.1 -100.5 0.49
1025 0.75 3.18 0.17 56.5 2.03 0.12 59.2 0.07 0.09 60.8 55.0 21.6 -92.9 0.31
Implementation Details. On the AF-200K dataset, we trained for 2 million iterations with a batch size of 2,048 using 8 NVIDIA 4090 GPUs. On the Supercritical Airfoil dataset, we trained for 1 million iterations with a batch size of 256 on a single NVIDIA 4090 GPU. We use the Matérn kernel function as the kernel for the Gaussian process at time $t=0$. Other training hyperparameters are detailed in the Supplementary Materials.
4.2 Main Results
Conditional Airfoil Generation.
As shown in Table 1, FuncGenFoil significantly outperforms the strongest baseline method (PK-DIT) across all metrics. For label error, FuncGenFoil achieves a reduction in average error by 88.5% on UIUC, 99% on Super, and 74.4% on AF-200K. This substantial decrease in label error underscores the effectiveness of our approach in generating airfoils that more precisely adhere to target geometric parameters. In terms of diversity, FuncGenFoil demonstrates notable improvements, surpassing the best baseline methods by 75.4, 14.6, and 21.6 on UIUC, Super, and AF-200K, respectively. This highlights the model’s superior capability in capturing and generating a broader spectrum of valid airfoil designs. Additionally, our generated airfoils exhibit enhanced surface smoothness, as evidenced by reductions of 0.99 and 4.86 in smoothness values ($10^{-2}$) on the UIUC and Super datasets, respectively. This improvement is particularly crucial for aerodynamic performance, as smoother airfoil surfaces contribute to reduced drag and improved flow characteristics.
@@ 4.2 Main Results
Freestyle Airfoil Editing. Table 2 shows the average performance across 300 editing cases using FuncGenFoil. In each case, the model adjusts 2 to 4 randomly selected positions on the wing surface to target locations over 10 fine-tuning steps, with edit scales ranging from $1\times 10^{-4}$ to $4\times 10^{-4}$. The results demonstrate that FuncGenFoil can perform accurate airfoil editing with minimal errors (less than $2.75\times 10^{-7}$ MSE) and high surface smoothness (less than 1.16 ×$\times 10^{-2}$ smoothness value) with just a few fine-tuning steps. Fig. 5 illustrates the editing requirements and generated airfoils during the fine-tuning stage. It shows that our model faithfully completes the freestyle editing task by generating reasonable airfoils according to user-specified editing requirements.
@@
Figure 5: Examples of two FuncGenFoil instances performing airfoil editing over 100 training iterations. The red section represents the editing requirement, while the black airfoil shows the generated samples at each training stage. The results demonstrate that the generated airfoil quickly adapts to the editing requirements within a few iterations, achieving a natural and smooth function regression.
@@
Any-Resolution Airfoil Generation. One of the key advantages of FuncGenFoil is its ability to generate airfoils at any resolution while maintaining high generation quality. This is achieved by representing airfoils as functions and learning resolution-independent function transformations, enabling flexible and consistent airfoil generation across different scales. To evaluate high-resolution generation, we use the model trained on the Supercritical Airfoil dataset at a resolution of 257 and sample new airfoils at resolutions of 513 and 1025. We then assess these higher-resolution airfoils using the same metrics as in the conditional generation task. The results are presented in Table 3. We observe consistent generation quality at $\times$ and even $\times$ the training resolution, with a maximum increase of 9.1×$\times 10^{-3}$ in average label error, a decrease of 16.1 in diversity, and an increase of 0.67×$\times 10^{-2}$ in smoothness value.
Aerodynamic Simulation. To further assess the physical properties of the generated airfoils and validate the effectiveness of our method, we analyze their aerodynamic performance using the NASA Common Research Model (CRM)  [^2] https://commonresearchmodel.larc.nasa.gov/ and perform Reynolds-Averaged Navier-Stokes (RANS) computational fluid dynamics (CFD) simulations on the generated samples. The CRM dataset contains 135,000 CRM wing geometries along with their corresponding aerodynamic performance, computed using the RANS CFD solver ADflow Mader et al. (2020). We pretrain the FuncGenFoil model on all 135,000 CRM wing geometries and generate 500 new CRM wing geometries for RANS CFD evaluation. We analyze the lift-to-drag ratio ($L/D$) of these newly generated geometries and compare them with the original dataset, as shown in Fig. 6. Our results indicate that the $L/D$ distribution of the generated samples closely aligns with that of the original dataset, with the highest density occurring around $L/D=19$, demonstrating the model’s ability to learn and generate physically plausible wing geometries. Additionally, we visualize the coefficient of pressure contours for selected CFD cases in Fig. 7. These visualizations confirm that FuncGenFoil can generate airfoils with diverse aerodynamic performance characteristics.
@@
Figure 6: Comparison of the lift-to-drag ratio histograms for CRM wings in the original dataset and the samples generated by our FuncGenFoil model.
Figure 7: Visualization of the pressure coefficient for the generated CRM wings, obtained through aerodynamic simulation.

### v3

@@ 3.2 Airfoil Editing
The airfoil editing task enables the user to modify parts of the geometry of a given airfoil $u_{1}$, effectively generating a new airfoil geometry $u_{1}^{{}^{\prime}}$ while preserving the user-edited sections, denoted as $u_{1}^{\delta}$. We achieve this by finetuning the pretrained model via maximum a posteriori (MAP) estimation, $\max p(u_{1}^{{}^{\prime}}\mid u_{1}^{\delta})$ where $p(u_{1}^{{}^{\prime}}\mid u_{1}^{\delta})$ is a probabilistic model that constrains the optimized airfoil $u_{1}^{{}^{\prime}}$ fulfilling the editing requirements and following the generation prior.
@@
$\max_{u_{1}^{{}^{\prime}}}p(u_{1}^{{}^{\prime}}\mid u_{1}^{\delta})\;=\;\frac{p(u_{1}^{\delta}\mid u_{1}^{{}^{\prime}})\cdot p(u_{1}^{{}^{\prime}})}{p(u_{1}^{\delta})}\Rightarrow\max_{u_{1}^{{}^{\prime}}}\log{p(u_{1}^{\delta}|u_{1}^{{}^{\prime}})}+\log{p(u_{1}^{{}^{\prime}})}-\log{p(u_{1}^{\delta})},$ (7)
where $p(u_{1}^{\delta}\mid u_{1}^{{}^{\prime}})$ is a Gaussian measure, so its log term becomes a Mean Square Error (MSE) between $u_{1}^{\delta}$ and $u_{1}^{{}^{\prime}}$, constraining the user edited parts in $u_{1}^{{}^{\prime}}$ keeping consistent with $u_{1}^{\delta}$. We can sample $u_{1}^{{}^{\prime}}$ through solving the velocity operator using Neural ODE method[6] and calculate $\log p(u_{1}^{{}^{\prime}})$ at the same time using FFJORD method [17] with Hutchinson trace estimator [22]. $p(u_{1}^{{}^{\prime}})$ is the prior supported by the trained generative model. $p(u_{1}^{\delta})$ is the marginal likelihood, which does not depend on $u_{1}^{{}^{\prime}}$. The final optimization target could be written as:
@@
where $\sigma$ is noise scale for editing, $\Delta$ denotes point indices in edited part $u_{1}^{\delta}$.
For more realistic generation results, the optimization does not directly adjust $u_{1}^{{}^{\prime}}$; instead, we optimize $u_{1}^{{}^{\prime}}$ indirectly by fine-tuning the entire generative model for a few iterations. The fine-tuning process is illustrated in Figure 3, while its details are provided in Alg. 3 in the Appendix. Specifically, we first initialize $u^{\prime}_{1}$ as a resample data of $u_{1}$, by extracting its latent code, denoted as $\alpha_{\theta}$, via the inverse of our generative model and re-generating $u_{1}^{{}^{\prime}}$ from this code. Then we treat Equation 8 as the loss function to train $\theta$ of the velocity operator and $\alpha_{\theta}$ simultaneously. After the model training, the new $u_{1}^{{}^{\prime}}$ is the new edited generation results.
@@ 4.1 Experimental Settings
Tasks. We evaluate two tasks within the airfoil inverse design problem: conditional airfoil generation and freestyle airfoil editing. In the conditional generation task, the model is provided with a set of 11 geometric parameters describing the airfoil geometry. Detailed parameter definitions are provided in Table 6 in the Appendix. The model must generate airfoils that satisfy these geometric constraints. In the freestyle editing task, the model receives an original airfoil and a target modification, such as adjusting the position of a specific point on the airfoil. The selected point can be located anywhere on the airfoil surface. The model must generate an airfoil reflecting the specified modification.
Metrics. We adopt the metrics introduced in AFBench [37] to evaluate the generated airfoils:
@@
Label Error measures the difference between the PARSEC parameters of the generated or edited airfoil and the intended target parameters, calculated as $\sigma_{i}=\left|\hat{p}_{i}-p_{i}\right|$, where $\sigma_{i}$ is the label error for the $i$-th parameter, $\hat{p}_{i}$ is the $i$-th geometric parameter of the generated airfoil, and $p_{i}$ is the corresponding target parameter. Smaller values indicate better alignment with target parameters. To summarize all $11$ label errors, we present both the arithmetic mean $\bar{\sigma}_{a}$ and the geometric mean $\bar{\sigma}_{g}$ for absolute and relative average errors.
@@
Diversity quantifies the variety among generated airfoils, calculated as $D=\frac{1}{n}\sum_{i=1}^{n}\log\det(L_{S_{i}})$, where $n$ is the number of subsets, and $L_{S_{i}}$ is the similarity matrix of the $i$-th subset, computed based on Euclidean distances between airfoils within the subset. Higher values indicate greater diversity among generated airfoils.
Smoothness measures the geometric smoothness of generated airfoils, calculated as:
$\small M=\sum_{n=1}^{N}\text{Distance}(P_{n},P_{n-1}P_{n+1}),$ (9)
@@
where $P_{n}$ is the $n$-th point, and $P_{n-1}P_{n+1}$ is the line segment connecting its adjacent points. The function $\text{Distance}(P_{n},P_{n-1}P_{n+1})$ computes the perpendicular distance from $P_{n}$ to this line segment. Smaller values indicate better geometric quality.
@@
Datasets. To benchmark our method, we conduct experiments on three datasets: UIUC [50], Supercritical Airfoil (Super), and AF-200K. UIUC contains 1,600 designed airfoil geometries. Super focuses on supercritical airfoils and includes approximately 20,000 airfoil samples. AF-200K consists of 200,000 highly diversified airfoil samples.
@@
Baselines. We compare FuncGenFoil with baseline models proposed in AFBench [37], specifically the conditional VAE (CVAE), conditional GAN (CGAN), the modified VAE with PARSEC parameters and control keypoints (PK-VAE), as well as PK-GAN, PKVAE-GAN, the U-Net-based PK-DIFF, and the transformer-based PK-DIT.
@@
Table 1: Quantitative comparison between FuncGenFoil and baseline methods across different datasets for the conditional generation task. Label error, diversity, and smoothness of generated airfoils are reported.
@@
$\sigma_{1}$ $\sigma_{2}$ $\sigma_{3}$ $\sigma_{4}$ $\sigma_{5}$ $\sigma_{6}$ $\sigma_{7}$ $\sigma_{8}$ $\sigma_{9}$ $\sigma_{10}$ $\sigma_{11}$ $\bar{\sigma}_{a}$ $\bar{\sigma}_{g}$
CVAE AF-200K 72.9 52.5 35.2 15900 99 95 29000 19.1 15.3 46 104 4131 149.2 -155.4 7.09
CGAN AF-200K 107 85.0 54.4 23200 143 137 59600 25.3 22.3 53 129 7596 217.6 -120.5 7.31
PK-VAE AF-200K 63.0 47.9 31.3 8620 66 64 17100 13.5 9.3 33 78 2375 106.0 -150.1 5.93
PK-GAN AF-200K 81.8 63.0 47.0 21030 120 117 32470 22.5 19.6 50 122 4922 179.5 -112.3 3.98
PKVAE-GAN AF-200K 56.8 31.7 31.0 5650 46 43 12000 9.1 5.1 28 63 1633 77.6 -129.6 2.89
PK-DIT AF-200K 11.2 32.3 15.4 1050 13 11.5 9790 0.5 0.5 23 24 997 23.5 -93.2 1.04
FuncGenFoil AF-200K 1.84 17.4 0.56 721 47.7 0.98 1676 0.45 0.65 160 174 255 14.9 -71.6 1.41
PK-DIT UIUC 63.8 51.4 33.6 11830 87 84.9 25700 16.9 11.9 36 98 3456 129.8 -141.7 6.03
FuncGenFoil UIUC 11.6 11.8 0.38 698 23.5 0.56 12339 0.27 0.44 110.5 115.0 1210 15.1 -91.1 1.33
PK-DIT Super 52.0 35.0 24.0 3010 29 33.2 10500 8.3 2.6 27 33 1250 58.1 -123.4 1.97
FuncGenFoil Super 0.71 8.23 0.13 201.3 4.72 0.12 174.2 0.09 0.14 34.2 36.7 41.9 3.08 -103.9 1.01
Table 2: Quantitative evaluation of the FuncGenFoil model across different sampling super-resolutions for the conditional generation task. The training resolution is $257$.
[…]
$\sigma_{1}$ $\sigma_{2}$ $\sigma_{3}$ $\sigma_{4}$ $\sigma_{5}$ $\sigma_{6}$ $\sigma_{7}$ $\sigma_{8}$ $\sigma_{9}$ $\sigma_{10}$ $\sigma_{11}$ $\bar{\sigma}_{a}$ $\bar{\sigma}_{g}$
Super 257 0.71 8.23 0.13 201.3 4.72 0.12 174.2 0.09 0.14 34.2 36.7 41.9 3.08 -103.9 1.01
513 0.71 8.19 0.12 200.8 4.81 0.12 174.7 0.09 0.15 48.2 47.8 44.1 3.26 -97.5 0.52
1025 0.71 8.29 0.12 202.9 4.93 0.12 175.4 0.09 0.15 61.4 57.8 46.5 3.44 -91.1 0.37
Implementation Details. On the AF-200K dataset, we trained for 2 million iterations with a batch size of 2,048 using 8 NVIDIA 4090 GPUs. On the Supercritical Airfoil and the UIUC dataset, we trained for 1 million iterations with a batch size of 1,024 on 4 single NVIDIA 4090 GPUs. We use Gaussian processes with a Matérn kernel function ($\nu=2.5,l=0.03$) as the prior. Other training hyperparameters are detailed in Appendix C.
4.2 Main Results
Conditional Airfoil Generation. As shown in Table 1, FuncGenFoil outperforms the strongest baseline method (PK-DIT) across all metrics. For label error, FuncGenFoil achieves reductions in arithmetic mean label error of 74.4% on AF-200K, 65.0% on UIUC, and 96.6% on Super. More importantly, it also reduces the geometric mean label error by 36.6% on AF-200K, 88.4% on UIUC, and 94.7% on Super. This substantial decrease in label error underscores the effectiveness of our approach for generating airfoils that more precisely adhere to target geometric parameters. In terms of diversity, FuncGenFoil demonstrates notable improvements, surpassing the best baseline methods by 23.2%, 35.7%, and 15.8% on AF-200K, UIUC, and Super, respectively. This highlights the model’s superior capability in capturing and generating a broader spectrum of valid airfoil designs. Additionally, the generated airfoils exhibit enhanced surface smoothness, as evidenced by reductions of 4.70 and 0.96 in smoothness values ($10^{-2}$) on the UIUC and Super datasets, respectively. This improvement is particularly crucial for aerodynamic performance, as smoother airfoil surfaces contribute to reduced drag and improved flow characteristics.
Table 3: Quantitative evaluation of the airfoil editing task across different editing scales. Mean squared error (MSE) between generated and target airfoils and smoothness of generated airfoils are reported.
@@ 4.2 Main Results
Freestyle Airfoil Editing. Table 3 shows average performance over 300 editing cases using FuncGenFoil. In each case, the model adjusts 2 to 4 randomly selected positions on the airfoil surface to target locations over 10 fine-tuning steps, with editing scales ranging from $1\times 10^{-4}$ to $4\times 10^{-4}$. Results demonstrate that FuncGenFoil achieves accurate airfoil editing with minimal errors (less than $2.75\times 10^{-7}$ MSE) and high surface smoothness (less than 1.16 ×$\times 10^{-2}$ smoothness value) after only a few fine-tuning steps. Figure 4 illustrates example editing requirements and corresponding generated airfoils during the fine-tuning stage, demonstrating that our model effectively completes freestyle editing tasks by generating accurate airfoils according to user-specified edits.
@@
Figure 4: Examples of eight different FuncGenFoil instances performing airfoil editing over 20 training iterations. The orange points and sections represent the editing requirements as constraints; the gray region is the original airfoil, while the blue region shows the generated edited airfoil. The results demonstrate that the generated airfoil quickly adapts to the editing requirements within a few iterations, achieving a natural and smooth function regression. The editing scheme can be completely customized according to the user’s preference.
@@
Any-Resolution Airfoil Generation. One of the key advantages of FuncGenFoil is its ability to generate airfoils at any resolution while maintaining high generation quality. This is achieved by representing airfoils as functions and learning resolution-independent function transformations, enabling flexible and consistent airfoil generation across different scales. To evaluate high-resolution generation, we use the model trained on the Supercritical Airfoil dataset at a resolution of 257 and sample new airfoils at resolutions of 513 and 1025. We then assess these higher-resolution airfoils using the same metrics as in the conditional generation task. The results are presented in Table 2. We observe consistent generation quality at $\times$ and even $\times$ the training resolution, with a maximum increase of 4.6×$\times 10^{-3}$ in average label error, a decrease of 12.8 in diversity, and an increase of 0.64×$\times 10^{-2}$ in smoothness value.
@@
Figure 5: Left: Comparison of the lift-to-drag ratio histograms for CRM wings in the original dataset and the samples generated by our FuncGenFoil model. Right: Visualization of the pressure coefficient for the generated CRM wings, obtained through aerodynamic simulation.
Aerodynamic Simulation. To further assess the physical properties of the generated airfoils and validate the effectiveness of our method, we analyze their aerodynamic performance using the NASA Common Research Model (CRM)  [^2] https://commonresearchmodel.larc.nasa.gov/ and perform Reynolds-Averaged Navier-Stokes (RANS) computational fluid dynamics (CFD) simulations on the generated samples. The CRM dataset contains 135,000 CRM wing geometries along with their corresponding aerodynamic performance, computed using the RANS CFD solver ADflow [39]. We pretrain the FuncGenFoil model on all 135,000 CRM wing geometries and generate 500 new CRM wing geometries for RANS CFD evaluation. We analyze the lift-to-drag ratio ($L/D$) of these newly generated geometries and compare them with the original dataset, as shown in Figure 5. Our results indicate that the $L/D$ distribution of the generated samples closely aligns with that of the original dataset, with the highest density occurring around $L/D=19$, demonstrating the model’s ability to learn and generate physically plausible wing geometries. Additionally, we visualize the coefficient of pressure contours for selected CFD cases in Figure 5. These visualizations confirm that FuncGenFoil can generate airfoils with diverse aerodynamic performance characteristics.